Solve for the unknown: 8/52 = m/39
_______________________________________________________________________________
Is the rate 1.2 inches of rain/5hours equivalent to the rate 2.1 inches of rain/8hours?

Answer by stanbon(48526)

Solve for the unknown: 8/52 = m/39
Multiply both sides by 39 to get:
m= 39(8/52) = 13(8/14)= 13(4/7) = 52/7

_______________________________________________________________________________
Is the rate 1.2 inches of rain/5hours equivalent to the rate 2.1 inches of rain/8hours?
1.2/5 ? 2.1/8
------------------
To get the 8 in the denominator multiply numerator and denominator
by (8/5).
You would get [(1.2*(8/5))]/[5*(8/5)) = 1.92/8
----------------
So, they are not equivalent.
=====================
Cheers,
Stan H.